OpenStudy (abmon98):

solve the following equation using the substitution 2^x=u 2^x+2^2-x=5 (4^x)-3(2^x)=10 2^(2x+1)-11(2^x)+15=0

OpenStudy (abmon98):

nope 2^x

OpenStudy (abmon98):

take your time these are three seperate equation

OpenStudy (anonymous):

We need to say somehow that u^2+something-u=5 so we can solve for u and then substitute 2^x in the answer to get our solution...

OpenStudy (anonymous):

Square each term in the equation maybe.

Parth (parthkohli):

\[2^x + 2^{2-x}=5\]Write it something like this:\[2^x + \dfrac{2^2}{2^x}=5\]It should be clearer now.

ganeshie8 (ganeshie8):

@Abmon98

ganeshie8 (ganeshie8):

just plug in 2^x = u and you will see a quadratic in u